osmium os is the densest element known density 22.57 cm3 calculate the mass in grams of a cylinder of osmium with a height of 3.0cm and a diameter of 1.0cm?
1
Expert's answer
2018-03-14T08:57:04-0400
Solution
m = Vcyl×ρ
Vcyl = πR^2 * H
m = 3.14 × 0.5^2 × 3 × 22.57 = 53.15 (g)

Answer
The mass of a cylinder is 53.15 grams.